When given the problem, x + 5y = 15, How do I solve for y?
Answer by sdmmadam@yahoo.com(530)

When given the problem, x + 5y = 15, How do I solve for y?
You are given a linear equation in x and y.
You can express y as a function in x as follows:
x + 5y = 15 ----(1)
5y = -x+15
(taking x to the other side, the rule being change side then change sign.)
Dividing by 5
y = (-1/5)x +3 ----(*)
Note:A linear equation in x and y reperesents the equation to a straight line and the given equation: x + 5y-15=0 called the equation to a straight line in the general form is converted into the form (*) called the slope and the y-intercept form of equation to the line.
When the coefficient of y on one side of the equation is 1, the coefficient of x on the other side gives the slope and the free constant on the other side gives the y-intercept.
slope here is negative implies the straight line is making an obtuse angle with the x -axis and hence its journey is from the II quadrant to the fourth quadrant and since the y-intercept is positive,the journey is via the first quadrant.The line meets the y-axis at (0,3) at a point 3 units above the origin.
Note: If you need not points on a whole line but one single point (x,y) then we must be furnished with one more linear equation in x and y such that the values for x and y hold in both the equations!